# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-Edge AI Applications

### v3

@@ 5.3. Cross-Hardware Support
To show the effectiveness of MSCCL++ on different hardware platforms, Figure 11 and 12 compare the single-node AllReduce performance on H100 and MI300x.
@@
MI300x. MSCCL++ outperforms RCCL and MSCCL by up to 3.8x and 1.9x for small messages, and by up to 2.2x and 1.6x for large messages, respectively. Our algorithm implementations take special considerations to best utilize the Infinity Fabric that peer-to-peer connects all GPUs in a node, unlike NVLink that connects all GPUs to a switch. This means, for best link utilization, we need to copy data to all peers at the same time as much as possible, unlike for NVIDIA GPUs where we can copy data to each peer one-by-one back-to-back. Using the MSCCL++ DSL, this is as easy as changing the order of two nested for loops, one that walks through data elements and the other that walks through different peers.
@@ 6. Related Work
NVIDIA Collective Communications Library (NCCL) (NVIDIA, 2025a) and AMD ROCm Collective Communications Library (AMD, 2025) (RCCL) are vendor supplied libraries for NVIDIA and AMD GPUs respectively. Cowan et al. (Cowan et al., 2022) improved over these libraries by allowing execution of collective algorithms specialized for a size and topology. In this paper, we showed that primitives provided by MSCCL++ performs better than the primitives of these libraries.
The primitive interface of MSCCL++ has similarities with some other works in terms of being a GPU-side API for communication. For example, the NVIDIA OpenSHMEM (NVSHMEM) (NVIDIA, 2025b) is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective functions such as nvshmemx_broadcastmem_warp. However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and quantitative comparison is difficult as NVSHMEM is not open source. ARK (Hwang et al., 2023b) is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
UCX, an open-source framework developed by Shamis et al. (Shamis et al., 2015) and libraries implementing the MPI standard (Thakur et al., 2005) offer a comprehensive set of APIs for high-performance computing (HPC) environments. In contrast, MSCCL++ is more specific to GPU acceleration.
@@
SCCL (Cai et al., 2021), TACCL (Shah et al., 2023), and TE-CCL (Liu et al., 2024) aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath. The MSCCL++ primitives can be used with the generated algorithms to capture their benefit.
@@
Finally, LLM inference frameworks like vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2025e) implement custom AllReduce kernels for achieving high performance communication. However, the custom implementations are not general-purpose and often only limited to single-node collective communication. MSCCL++, on the other hand, enables communication across multiple types of interconnects and has equivalent AllReduce performance to vLLM’s custom single-node AllReduce.
7. Conclusion
@@
MSCCL++ is a novel GPU communication stack designed for high-performance AI applications. By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enable fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations. Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies. By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to $5.4\times$ speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.

### v4

@@ 7.2. Cross-Hardware Support
To show the effectiveness of MSCCL++ on different hardware platforms, Figure 9 and 10 compare the single-node AllReduce performance on H100 and MI300x.
@@
MI300x. MSCCL++ outperforms RCCL and MSCCL by up to 3.8x and 1.9x for small messages, and by up to 2.2x and 1.6x for large messages, respectively. Our algorithm implementations take special considerations to best utilize the Infinity Fabric that peer-to-peer connects all GPUs in a node, unlike NVLink that connects all GPUs to a switch. This means, for best link utilization, we need to copy data to all peers at the same time as much as possible, unlike for NVIDIA GPUs where we can copy data to each peer one-by-one back-to-back. Using the MSCCL++ DSL, this is as easy as changing the order of two nested for loops, one that walks through data elements and the other that walks through peers.
Figure 11. Speedup of MSCCL++ over NCCL for Llama3-70b decodes with tensor parallelism = 8.
7.3. LLM Inference
[…]
vLLM also has a custom hand-written AllReduce kernel for a single node. For different message sizes, MSCCL++’s AllReduce performs similar or up to $3\times$ faster than the custom kernel, with a geomean improvement of $1.4\times$. For end-to-end decode latency, MSCCL++ performs similar or up to $1.11\times$ (on average $1.04\times$) faster than inference with the custom AllReduce kernel.
[…]
Figure 12. DeepSeek-V3 decode throughput using SGLang on two H100 nodes (16 GPUs) with tensor parallelism (1024 input tokens and 1024 output tokens per batch).
Figure 13. DeepEP on two H100 nodes (16 GPUs) with expert parallelism (DeepSeek-V3 settings: hidden size 7168, top-k 8, 256 total experts, FP8 dispatch and BF16 combine).
Expert Parallelism with DeepEP. DeepEP (Zhao et al., 2025) is an implementation of expert parallelism for Mixture-of-Experts (MoE) layers, which is proposed by DeepSeek-V3 (DeepSeek-AI, 2024) for both training and inference. The original code leverages a hardware feature called InfiniBand GPUDirect Async (IBGDA) (Markthub et al., 2022) (featured by NVSHMEM (NVIDIA, 2025e)) that implements an InfiniBand networking stack inside GPU, so that GPU can directly communicate with NIC bypassing CPU. However, this feature is only available on NVIDIA GPUs combined with Mellanox NICs at the time of writing. To make it portable to other environments, we integrate MSCCL++ into DeepEP to replace the IBGDA-based networking functions into PortChannel API calls, which uses the RDMA networking stack running on CPU instead (see Figure 4). [^4] IBGDA can be supported by the MSCCL++ PortChannel API in the future. It took a couple of weeks for one developer to add this support, but most of that time was spent in understanding DeepEP code. We evaluated DeepEP using its official benchmark and setup. IBGDA was enabled for all experiments, which was leveraged by NVSHMEM, but not by MSCCL++. Figure 13 demonstrates that the two implementations exhibit no noticeable performance difference. Unlike NCCL, MSCCL++ is flexible enough to implement DeepEP-like applications efficiently. The IBGDA stack implementation in DeepEP is less portable and hard to maintain, while MSCCL++’s PortChannel simplifies this code drastically while being portable and performant.
7.4. Portability and Implementation Effort
[…]
The MSCCL++ design makes it easy to extend the Primitive API to support new hardware features. For example, NVIDIA NVLink 4.0 introduced support for multimem instructions for switch-based aggregation and multicast. Since this feature introduced a very different concept from previous interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature. The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than $2.2\times$ on average for message sizes of 1KB - 1GB.
@@ 8. Related Work
GPU Communication Primitives: NVSHMEM (NVIDIA, 2025e) is an extension of the OpenSHMEM (OpenSHMEM, 2025) API for NVIDIA GPUs for providing a shared-memory abstraction. In contrast, MSCCL++ provides performance-preserving portable abstractions and higher-level DSL for developing efficient communication collective algorithms. Concurrent to our work, NVSHMEM recently introduced support for inter-node communication (after MSCCL++ was open sourced). MSCCL++ provides different interconnect abstractions, namely PortChannel, MemoryChannel, and SwitchChannel, to enable optimized implementations. As an example, NVSHMEM’s put always uses thread-copy when it is feasible, so users cannot choose DMA-copy on NVLink. Similarly, NVSHMEM exposes raw multimem capabilities by exposing underlying memory pointers and users need to implement their kernels using PTX instructions. With an eye towards portability, SwitchChannel abstracts low-level details with a uniform synchronization mechanisms. Finally, NVSHMEM does not provide an API like MemoryChannel’s LL protocol. It may be possible to implement such an API using NVSHMEM, but that would add burden on users.
ARK (Hwang et al., 2023b) is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
DSL: Cowan et al. (Cowan et al., 2022) introduced MSCCLang, a DSL to specify collective communication algorithms specialized for a message size and network topology. MSCCLang is built on top of NCCL and RCCL send-recv abstraction, so is limited to two-sided, synchronous communication. Our MSCCL++ DSL provides a one-sided, asynchronous API that enables expressing more efficient communication algorithms. Due to this, MSCCL++ is faster than MSCCL, which is built using MSCCLang DSL, as shown in our evaluation.
@@
Communication Libraries: NVIDIA Collective Communications Library (NCCL) (NVIDIA, 2025d) and AMD ROCm Collective Communications Library (AMD, 2025) (RCCL) are vendor supplied libraries for NVIDIA and AMD GPUs respectively. We showed that MSCCL++ Collectives perform better than both for AI inference workloads. RCCL has since adopted MSCCL++.
@@
Communication Algorithms: SCCL (Cai et al., 2021), TACCL (Shah et al., 2023), and TE-CCL (Liu et al., 2024) aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms using the NCCL and RCCL send-recv abstraction. Their techniques are orthogonal and their performance may improve if they generate algorithms using the MSCCL++ Primitives API.
@@
LLM Inference Frameworks: LLM inference frameworks like SGLang (SGLang, 2025), DeepEP (Zhao et al., 2025), vLLM (Kwon et al., 2023), and TensorRT-LLM (NVIDIA, 2025g) implement custom kernels for achieving high performance communication. However, the custom implementations are not general-purpose and often only limited to specific interconnects like NVLink or xGMI. MSCCL++ not only performs similar to their custom implementations on their corresponding interconnects, but also supports efficient communication across multiple types of interconnects. SGLang has since adopted MSCCL++.
9. Conclusion and Future Work
@@
MSCCL++is a novel GPU communication stack designed for high-performance AI applications. By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enables fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations. Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies. By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to $5.4\times$ speedup for standalone collectives, and up to $1.38\times$ speedup for end-to-end AI inference. MSCCL++ could be used for accelerating AI training workloads too, but we leave that for future work. MSCCL++ could also evolve into a general I/O abstraction for AI accelerators in the future; e.g., by supporting non-GPU accelerators and file system I/O transports like BlitzScale (Zhang et al., 2025).
Appendix A Artifact Appendix
A.1. Abstract
The artifact includes implementations of the three APIs described in this paper, namely the Primitive API (C++/CUDA) for low-level optimization of GPU communication, the DSL API (Python) that is implemented over the Primitive API and enables high-level description of custom collective communication algorithms, and the Collective API (C++) that implements the NCCL/RCCL API as-is over the Primitive API and the DSL API. The Collective API can be used to reproduce collective communication results in Section 7.1 and 7.2.
A.2. Artifact check-list (meta-information)
  • Program: The latest version of nccl-tests (or rccl-tests). Publicly available and not included.
  • Run-time environment: Linux; Ubuntu 20.04 or later is recommended. Docker is recommended, otherwise need manual installations of dependencies such as CUDA (or ROCm), Python3, MPI, and optionally OFED for multi-node experiments.
  • Hardware: 8 datacenter-grade GPUs. Optionally RDMA NICs for multi-node experiments.
  • Execution: Sole user recommended for best performance.
  • Metrics: Latency and bandwidth are reported.
  • Output: Console output with reported metrics.
  • Experiments: Step-by-step guide in documentation.
  • How much disk space required (approximately)?: ¡1GB for the source code and a build; 10GB for a Docker image with all dependencies.
  • How much time is needed to prepare workflow (approximately)?: 10 minutes
  • How much time is needed to complete experiments (approximately)?: 1 hour
  • Publicly available?: Yes
  • Code licenses (if publicly available)?: MIT License
  • Archived (provide DOI)?: 10.5281/zenodo.18334822
A.3. Description
A.3.1. How to access
The source code is publicly available on GitHub at https://github.com/microsoft/mscclpp. The repository is sized less than 1MB, and a build would be a few hundreds of MB.
A.3.2. Hardware dependencies
At least 2 GPUs are needed for functionality tests, and 8 or more are needed for reproducing numbers in Section 7.1 and 7.2. Datacenter-grade GPUs with high-bandwidth interconnects (NVLink or Infinity Fabric) are recommended (see Table 2). Optionally for multi-node experiments, RDMA NICs with InfiniBand or RoCE configurations are required.
A.3.3. Software dependencies
CUDA (or ROCm), Python3, and MPI are required for artifact evaluation. Optionally OFED is required for multi-node experiments. See details in the ”Prerequisites” section in docs/quickstart.md.
A.4. Installation
  1. (1)
    Download and run a provided Docker image. This step can be skipped if the host environment is preferred. See ”Docker Images” section in docs/quickstart.md.
  2. (2)
    Build from the source. See the step-by-step guide in ”Install from Source” section in docs/quickstart.md. If the environment is not configured for multi-node execution, we recommend to disable InfiniBand/RoCE features via a CMake option -DMSCCLPP_USE_IB=OFF.
  3. (3)
    Build and run unit_tests and mp_unit_tests for functionality testing. See the step-by-step guide in ”Unit Tests” section in docs/quickstart.md.
A.5. Evaluation and expected results
For evaluation, install the latest version of nccl-tests (for NVIDIA GPUs) or rccl-tests (for AMD GPUs), which are publicly available at https://github.com/NVIDIA/nccl-tests and https://github.com/ROCm/rccl-tests, respectively. MPI should be enabled during their compilation (set environment variable MPI=1 during make).
After that, follow the ”NCCL/RCCL Benchmark over MSCCL++” section in docs/quickstart.md to run nccl-tests or rccl-tests using the MSCCL++ Collective API. The testing binaries will print latency and bandwidth numbers on the console, similar to what presented in Figure 7, 8, 9, and 10, depending on the GPU type and the number of GPUs. Please refer to the README.md of nccl-tests for detailed usages.
A.6. Experiment customization
The implementation of MSCCL++ Collective API can be customized using two different interfaces.
  • Using the Primitive API. Users can write their own GPU communication kernel using the MSCCL++ Primitive API and let the Collective API invoke the custom kernel. We provide a simple example in
    docs/guide/customized-algorithm-with-nccl-api.md.
  • Using the DSL API. Users can leverage our Python DSL to describe a collective communication algorithm that can be lowered and executed by the Collective API. We provide a simple example in docs/dsl_quick_start.md.
A.7. Notes
All documentation is also publicly available at https://microsoft.github.io/mscclpp.
A.8. Methodology
Submission, reviewing and badging methodology: